OpenStudy (anonymous):

Given the position function, s of t equals t cubed divided by 3 minus 12 times t squared divided by 2 plus 36 times t , between t = 0 and t = 15, where s is given in feet and t is measured in seconds, find the interval in seconds where the particle is moving to the right.

OpenStudy (astrophysics):

\[s(t) = \frac{ t^3 }{ 3-12t^2+36t }\] this?

OpenStudy (anonymous):

yes

OpenStudy (astrophysics):

wait no it's not..

OpenStudy (astrophysics):

\[s(t) = \frac{ t^3 }{ 3 }-\frac{ 12t^2 }{ 2 }+36t\]

OpenStudy (astrophysics):

Yeah ok there we go

OpenStudy (anonymous):

yeah sorry

OpenStudy (astrophysics):

Do you know calculus?

OpenStudy (astrophysics):

Finding the derivative will give you the velocity

OpenStudy (anonymous):

yes somewhat

OpenStudy (astrophysics):

Can you find the derivative of your function

OpenStudy (anonymous):

(x-6)^2?

OpenStudy (astrophysics):

How did you get that

OpenStudy (astrophysics):

Try the power rule, mhm do you know what derivatives are?